# Adversarial Markov Games: On Adaptive Decision-Based Attacks and Defenses

### v2

@@ IV Environment Specification
[Threat Model]. Our AMG framework describes a two-player competitive game; while extensible to more players, in this work we assume that at a given moment only one attack takes place. From the defensive perspective, incoming queries can be either benign or part of an attack. An assumption that influences the effectiveness of stateful detection is that queries can be attributed to UIDs, e.g., an IP address or a user account. However, adversaries can collude, create multiple accounts, use VPNs, or in fact accounts and IP addresses might not even be required to query the model. To address this, we treat queries irrespective to their source. This is a strictly more challenging setting for stateful defenses, where we operate solely on the content of queries and not on any other metadata, similar to [23]. Unlike Blacklight however, instead of rejecting queries, something that by itself can provide more information to the adversary and thus facilitate evasion (cf. OARS [24], LABEL:tab:comparison), we misdirect by returning the second highest probability class. Furthermore, Gaussian noise is added to the benign queries to simulate a noisy channel and a shift in distribution and so that is not trivial for a defense to tell adversarial noise apart. In summary, the black-box threat model we consider is delineated as follows:
@@
  • Adversary Knowledge: The model $\mathcal{M}$ is known as the black-box function that transforms inputs $x\in[0,1]^{d}$ to outputs $c\in[m]$, $m$ being the number of classes. The weights ${w}$ and the closed-form expression of $\mathcal{M}$ are unknown, as unknown is if an active defense $\pi_{\phi}^{\mathcal{D}}$ exists or not.
@@
  • Adversary Capabilities: Adapt the parameters of the attack and optionally any evasive transformations; in essence, optimize the candidate generation policy $\pi^{\mathcal{A}}_{\theta}$.
  • Defender Capabilities For each query $x$, decide between answering truthfully with the true prediction $C(x)=c_{T}$ or misdirect with the second highest probability class $c_{S}$.
Figure 3: Misdirection in a two-dimensional decision boundary. The adaptive defense controls a single parameter, the hypersphere radius around $k_{0}$ (the last known adversarial); for queries $x_{t}$ that fall within this hypersphere the model responds with a non-adversarial decision. $x_{g}$ is the starting sample, $x_{c}$ the original, and $x_{b}$ the best possible adversarial.
[Similarity]. Decision-based attacks typically follow a policy that generates successive queries: these exhibit degrees of similarity which can be quantified by an appropriate $l_{p}$ norm. If that norm is computed on the original inputs however, an adversary can adapt (see Figure 1) by employing evasive transformations the model is invariant to and bypass the similarity detection. To account for this capability, we train a Siamese network with contrastive loss in order to learn a latent space $\mathcal{L}(\cdot)$ where similar inputs are mapped close together, unaffected by added noise or transformations on the inputs. For the stateful characterization of queries, we use two queues: one for the detected adversarial queries as determined by the defensive agent, and one for the benign or undetected ones.
@@
[Adaptivity]. No evaluation in AML is complete without considering adaptive adversaries; a notion we expand in this work, that is with the instruments to bypass the defense and their optimal configuration. As stateful defenses are so far similarity based, to bypass them intuition points towards input transformations the model is invariant to. For a given query $x_{t}$ we want to compute a transformation $x_{t}^{\prime}=T(x_{t})$ so that $||x_{t}^{\prime}-x_{t}||_{2}\gg||x_{t}-x_{t-1}||_{2}$ while $F(T(x_{t}))\approx F(x_{t})$. Depending on magnitude and composition of transformations $T$, the identity $F(T(x_{t}))=F(x_{t})$ might not always hold. As we also demonstrate in Section V, $T$ interferes with the perturbations the adversarial policy generates itself: the performance and evasiveness of an attack are thus in a natural trade-off.
At this point we should inquire what is the correct composition of transformations $T$ to apply. When shall $T$ be applied, and how does it affect the attack fundamentals? The transformations $T$ can be considered as a set of additional controls, and like attack parameters they themselves can be suboptimal out-of-the-box [15]. Thus the combined control of attack and evasion parameters is a prerequisite to properly assess the strength of a defense. Their trade-off illustrates why the twofold definition of adaptive is necessary in AML evaluations: first to impart the tools to accomplish to the task through the definition of what can be controlled, and then to find the precise optimal configuration and strategy of the attack.
[Agents & Environments]. Unlike common competitive games, in AMGs the two players have different action and state sets. AMGs are also asymmetric in the playing cadence: while the defender plays every round, the adversary might wait one to several rounds; HSJA for example is controlled on the iteration rather on the query level. Training is complicated further given that the experience upon which each agent learns arrives only after the opponent moves. We address these complications by developing custom learning environments for agents with asymmetries, with delayed experience collection, and asynchronous training, build with the OpenAI Gym and Stable-Baselines3 libraries.
@@ V Evaluation
For evaluation, we define a range of scenarios intended to reflect all possible and realistic combinations between adversarial attacks and defenses, and their adaptive versions. Concretely, the research questions we want to validate are: 1) Are active defenses a necessary complement to model hardening and to what extent? 2) Are attacks more threatening when adaptive, i.e., do they outperform their vanilla versions and evade active detection? 3) If yes, can active defenses recoup their performance by also turning adaptive?
@@ V-A Evaluation Setup
The black-box attacks we render adaptive and evaluate are BAGS and HSJA, as they represent two fundamentally different approaches, are highly effective, and have the highest evasion potential [23]. BAGS is a stochastic, search-based method where every query submitted is a new and potentially better adversarial example. Contrastively, HSJA is deterministic and composed of 3 different stages where the queries are generated in an aggregated manner: the vast majority of them are not candidate adversarial examples but means to approximate the gradient at the decision boundary.
In training and evaluation, the adversarial game is played as follows: the adversary starts by submitting a query, then the defender responds either truthfully (the true model prediction) or by misdirecting (the second highest probability class). Then the environment decides with chance $p$ if the adversary moves next, otherwise a benign query is drawn. In either case, it is the defender’s turn; in testing they are also oblivious to the nature of the query and know only the content. All experiments are performed with $p=0.5$; we also evaluate our trained defense when no attack is present ($p=0$) in Appendix C.
The scenarios for all possible combinations of (non-) adaptive attacks and defenses are repeated over two different datasets – CIFAR10 and MNIST – and over two models with the same architecture but with different training regimes: one with adversarial training and one without. The transition from single-agent to multi-agent RL hides challenges: we approach the AMG as a belief-state MDP (the requirement of knowing the exact opponent policies is relaxed) and use PPO [42] agents to discover optimal policies that will also constitute best responses for the full AMG [47]. However, learning independently of other agency breaks the theoretical guarantees of convergence [49], eg. in scenarios 7 & 8 where both agents learn simultaneously. The complete list of scenarios is:
TABLE II: ASR and mean $L_{2}$ perturbation for 1K, 2K, and 5K queries for CIFAR-10, against normally and adversarially trained models. Clean Acc. reports the accuracy on benign queries of the base model plus any defenses present; in the first two scenarios (no active defense) the baseline clean accuracy is reported. Yellow scenarios denote the baseline attack performance, green denote active and/or adaptive defenses, and red denote adaptive attacks. The asterisk denotes where input transformations were used for evasion.
@@
    VA-ND – Vanilla Attack / No Defense: Baseline performance of attacks (BAGS & HSJA) out-of-the-box, without any active defense.
@@
    AA-ND – Adaptive Attack / No Defense: How more optimal is the adaptive version of an attack compared to the baseline.
@@
    AA-VD – Adaptive Attack / Vanilla Defense: Similar to scenario (2), but now the attack is adaptive.
@@
    AA-TD – Adaptive Attack / Trained Defense: After the adaptive defense is optimized, its policy is fixed and an adaptive attack is trained against it.
@@
    AA-AD – Adaptive Attack / Adaptive Defense: The first scenario where both agents learn simultaneously, making the environment non-stationary. In practice, the convergence will vary and depend on the chosen hyperparameters and rewards. Here we report the best-case for the attack.
@@
    AA-AD – Adaptive Attack / Adaptive Defense: The exact setup as scenario 7, but the best-case for the defense is reported instead.
For each successive scenario we evaluate with the most successful policy found, as this is best practice in Markov Games: the worst case opponent policy is fixed and then a counter to it is learned [31, 38]. Fixing other policies when computing a best response is representative of learning in environments that can contain multiple agents, with the additional benefit of converting the problem to single-agent that, as detailed in Section III-C, one can solve with RL.
Comparison to SOTA. In Scenarios 0-8 we evaluate all possible combinations between our attack and our defense. As a baseline to compare to, we additionally evaluate our approach to the state-of-the-art stateful defenses and adaptive attacks, that is Blacklight [23] and OARS [24] respectively. We implement both Blacklight and OARS in our interactive environments by using their publicly available code and parameters. As our environments do not return a rejection signal and to make a fair comparison, for OARS rejection coincides with a non-adversarial decision. We define 5 further scenarios:
@@
    VA-BD – Vanilla Attack / Blacklight Defense: Baseline performance of the attacks against Blacklight.
@@
    OA-BD – OARS Attack / Blacklight Defense: OARS against Blacklight.
@@
    AA-BD – Adaptive Attack / Blacklight Defense: Our adaptive attack against Blacklight.
@@
Our experiments were run on multiple machines, however to give an idea for the time complexity of our defense, on an Intel i7-7700 CPU one forward pass in CIFAR – that is one response to one query – takes $8\pm 1.4$ ms for 700 MFLOPs.
@@ V-B Results
For consistency and comparability between evaluations, all results are from the same 100 test episodes. The gap value denotes the $L_{2}$ perturbation that initially separates the starting and best adversarial example, averaged over the 100 episodes. By testing the trained agents on budgets higher than 5K we discovered that the trend in reducing $L_{2}$ holds; to make the agent training tractable and the evaluation wider however, we limit the maximum query budget per adversarial episode to 5K. Tables II & III report the results for CIFAR10, while IV reports MNIST. Overall, the empirical results help us extract and highlight several important insights, practical observations, and general implications for the broader AML field:
@@
  • Our adaptive defense (AD) outperforms both Blacklight (BD) and non-adaptive stateful (VD), also when transferred (Scen. 12). In HSJA, it reduces ASR by $\sim$90when trained against OA specifically, while it offers similar protection to BD when transferred from another attack.
  • Even against the strongest attacker and for the worst case for the defense (Scen. 7), our AD achieves an ASR $\leq 23\%$.
  • Our adaptive attack (AA) outperforms OARS (Scen. 11) and vanilla attacks (VA) by a wide margin, without access to rejection sampling. and irrespective of the defense it faces; the one exception is our adaptive defense (AD), against which it has very limited success.
  • Evasive transformations interfere with the attack operation, exemplified by the difference between BAGS and HSJA in Scen. 5; for the attack to reach its full potential, they have to be adaptively controlled together.
  • In Scen. 1-8, where agents train against the best opponent policy as previously discovered, ASR oscillates as following a fixed policy enables the learning of an optimal counter to it; and eventually plateaus, indicating an equilibrium.
  • The advantage of AA is much more pronounced against active defenses, where they significantly outperform non-adaptive versions.
  • The performance of both attacks deteriorates considerably against active defenses, however the defenses reach their full potential only when also adaptive.
  • The first time active defenses effectively resist adaptive attacks is in Scen. 5 of CIFAR; we employed evasive transformations from then onward.
  • Different attack fundamentals respond differently to defenses; the gradient estimation stage of HSJA is naturally disadvantaged against similarity detection, while the jump and binary stages are advantaged.
  • When devising the adaptive defense for HSJA, it proved nearly impossible to engineer a state the agent can learn on by leveraging our knowledge of the attack and its geometric functioning. What did prove effective, however, was pure computation [^2] This reminds us of Sutton’s Bitter Lesson [50], the observation that progress in AI is often driven by gains in computation rather than problem-specific expert knowledge.: we learn an embedding with Contrastive Learning [51] from raw input to use a state, which transfers exceedingly well to other attacks like BAGS.
@@ VI Discussion
Limitations. To keep the amount of evaluations practical, we narrowed the scope to targeted attacks and to $L_{2}$ as the more suitable norm for visual similarity. Targeted attacks are strictly more difficult to perform than untargeted, while for binary classification targeted and untargeted coincide; our framework, however, can accommodate any adversarial goal or metric. Another simplifying assumption we make is that only one attack can take place at a time; however, the queuing technique we use for incoming queries is readily extensible to handle concurrent attacks. While we demonstrate that our active defense does transfer between attacks, another possibility to explore is training the defense on multiple kinds of attacks. Finally, in our evaluation we focus on a wide range of adaptive and non-adaptive scenarios where agents learn and adapt interactively, thus limiting the number of datasets we experiment with; we back our empirical study however with an extensive theoretical analysis that supports the generality of our findings independent of context.
Future Work. The AMG framework we introduce is general by design and can accommodate the learning of optimal offensive and defensive policies in any domain of interest beyond image classification. A promising path for future research is the extension of our adaptive attacks and defenses to other domains and modalities, for instance malware, bot, and network intrusion detection. This is specifically because our approach circumvents the main obstacle of mapping gradient-based perturbations to feasible objects (eg. binaries) and instead can function directly in the problem space [36]. Another compelling and formidable challenge is automating the adaptive evaluations in AML, that is adapting beyond a specification by inventing instruments to bypass defenses and thus imparting controllability to adversarial tasks. Finally, in our work we considered opponent agency as part of the environment; other domains, like malware detection, might benefit from explicit opponent modeling.

### v3

@@ IV Threat Model
[Threat Model]. Our AMG framework describes a two-player competitive game; while extensible to more players, in this work we assume that at a given moment only one attack takes place. From the defensive perspective, incoming queries can be either benign or part of an attack. An assumption that influences the effectiveness of stateful detection is that queries can be attributed to UIDs, e.g., an IP address or a user account. However, adversaries can collude, create multiple accounts, use VPNs, or in fact accounts and IP addresses might not even be necessary to query the model. To address this, we treat queries irrespective to their source. This is a strictly more challenging setting for stateful defenses, where we operate solely on the content of queries and not on any other metadata, similar to [19]. Unlike Blacklight however, instead of rejecting queries, something that in itself provides more information to the adversary and thus facilitates evasion (cf. OARS [18], LABEL:tab:comparison), we misdirect by returning the second highest probability class. Furthermore, Gaussian noise is added to the benign queries to simulate a noisy channel and a shift in distribution, so that is not trivial for a defense to tell adversarial noise apart. In summary, the black-box threat model we consider is delineated as follows:
@@
  • Adversary Knowledge: The model $\mathcal{M}$ is known as the black-box function that transforms inputs $x\in[0,1]^{d}$ to outputs $c\in[m]$, $m$ being the number of classes. The weights ${w}$ and the closed-form expression of $\mathcal{M}$ are unknown, as unknown is if an active defense $\pi_{\phi}^{\mathcal{D}}$ is present or not.
@@
  • Adversary Capabilities: Adapt the parameters of the attack and of any evasive transformations; in essence, optimize the candidate generation policy $\pi^{\mathcal{A}}_{\theta}$.
  • Defender Capabilities: For each query $x$, decide between answering truthfully with the actual prediction $C(x)=c_{T}$, or misdirect with the second highest probability class $c_{S}$.
[Similarity]. Decision-based attacks typically follow a policy that generates successive queries: these exhibit degrees of similarity which can be quantified by an appropriate $L_{p}$ norm. If that norm is computed on the original inputs however, an adversary can adapt by employing evasive transformations the model is invariant to and bypass the similarity detection (cf. Figure 1). To account for this capability, we train a Siamese network with contrastive loss in order to learn a latent space $\mathcal{L}(\cdot)$ where similar inputs are mapped close together, unaffected by added noise or transformations on the inputs. For the stateful characterization of queries, we use two queues: one for the detected adversarial queries as determined by the defensive agent, and one for the benign and undetected ones.
@@
[Adaptivity]. No evaluation in AML is complete without considering adaptive adversaries; a notion we expand in this work, that is with the instruments to bypass the defense and their optimal configuration. As stateful defenses are so far similarity based, to bypass them intuition points towards input transformations the model is invariant to. For a given query $x_{t}$ we want to compute a transformation $x_{t}^{\prime}=T(x_{t})$ so that $||x_{t}^{\prime}-x_{t}||_{2}\gg||x_{t}-x_{t-1}||_{2}$ while $F(T(x_{t}))\approx F(x_{t})$. Depending on magnitude and composition of transformations $T$, the identity $F(T(x_{t}))=F(x_{t})$ might not always hold. As we also demonstrate in Section V, $T$ interferes with the perturbations of the adversarial policy: the performance and evasiveness of an attack are thus in a natural trade-off.
At this point one should inquire what is the correct composition of transformations $T$ to apply. When shall $T$ be applied, and how does it affect the attack fundamentals? The transformations $T$ can be considered as a set of additional controls, and like attack parameters they themselves can be suboptimal out-of-the-box [15]. Thus the combined control of attack and evasion parameters is a prerequisite to properly assess the strength of a defense. Their trade-off illustrates why the twofold definition of adaptive is necessary in AML evaluations: first to impart the tools to accomplish to the task through the definition of what can be controlled, and then to find the precise optimal configuration and strategy of the attack.
Figure 3: Misdirection in a hypothetical 2D decision boundary. The adaptive defense controls a single parameter, the hypersphere radius around $k_{0}$ (the last known adversarial); for queries $x_{t}$ that fall within this hypersphere the model responds with a non-adversarial decision. $x_{g}$ is the starting sample, $x_{c}$ the original, and $x_{b}$ the best possible adversarial.
[Agents & Environments]. Unlike common competitive games, in AMGs the two players have different action and state sets. AMGs are also asymmetric in the playing cadence: while the defender plays every round, the adversary might wait one to several rounds; HSJA for example is controlled on the iteration rather on the query level. Training is complicated further given that the experience upon which each agent learns arrives only after the opponent moves. We address these challenges by developing custom learning environments (with the OpenAI Gym and Stable-Baselines3 libraries) for asymmetrical agents, with delayed experience collection, and asynchronous training.
@@ V Evaluation
For evaluation, we define a range of scenarios intended to reflect all possible and realistic combinations between adversarial attacks and defenses, and their adaptive versions. Concretely, the research questions we want to evaluate are: 1) Are active defenses a necessary complement to model hardening and to what extent? 2) Are attacks more threatening when adaptive, i.e., do they outperform their vanilla versions and evade active detection? 3) If yes, to what extend active defenses recoup their performance by also turning adaptive?
@@ V-A Evaluation Setup
The black-box attacks we render adaptive and evaluate are BAGS and HSJA, as they represent two fundamentally different approaches, are highly effective, and have the highest evasion potential [19]. BAGS is a stochastic, search-based method where every query submitted is a new and potentially better adversarial example. Contrastively, HSJA is deterministic and composed of 3 different stages where the queries are generated in an aggregated manner: the vast majority of them are not candidate adversarial examples but means of approximating the gradient at the decision boundary.
In training and evaluation, the adversarial game is played as follows: the adversary starts by submitting a query, then the defender responds either truthfully (the actual model prediction) or by misdirecting (the second highest probability class). Then the environment decides with chance $p$ if the adversary moves next, otherwise a benign query is drawn. In either case, it is the defender’s turn; during testing they are also oblivious to the nature of the query and know only the content. All experiments are performed with $p=0.5$; we also evaluate our trained defense when no attack is present ($p=0$) in Appendix C.
TABLE II: ASR and mean $L_{2}$ perturbation for 1K, 2K, and 5K queries for CIFAR-10, against normally and adversarially trained models. Clean Acc. reports the accuracy on benign queries of the base model plus any defenses present; in the first two scenarios (no active defense) the baseline clean accuracy is reported. Yellow scenarios denote the baseline attack performance, while green and red denote defensive and offensive scenarios respectively. The asterisk denotes where input transformations were used for evasion.
@@
The scenarios for all possible combinations of (non-) adaptive attacks and defenses are repeated over two datasets – CIFAR10 and MNIST – and over two models with the same architecture but different training regimes: with and without adversarial training. As the transition from single to multi-agent RL introduces non-stationarity, we approach the AMG as a belief-state MDP (relaxing the requirement of knowing the exact opponent policies), and use PPO [43] agents to learn optimal policies that will also constitute best responses for the full game [48]. Note that learning independently of other agency breaks the theoretical guarantees of convergence [50], eg. in scenarios 7 & 8 where both agents learn simultaneously. Coloring denotes the learning/evaluated agent in each scenario, with their complete list being as follows:
@@
    VA-NDVanilla Attack / No Defense: Baseline performance of attacks (BAGS & HSJA) out-of-the-box, without any active defense.
@@
    AA-NDAdaptive Attack / No Defense: How much more optimal is the adaptive version compared to the baseline attack.
@@
    AA-VDAdaptive Attack / Vanilla Defense: Similar to scenario (2), but now the attack is adaptive.
@@
    AA-TDAdaptive Attack / Trained Defense: After the adaptive defense is optimized, its policy is fixed and an adaptive attack is trained against it.
@@
    AA-ADAdaptive Attack / Adaptive Defense: The first scenario where both agents learn simultaneously, making the environment non-stationary. In practice, the convergence will vary and depend on the chosen hyperparameters and rewards. Here we report the best-case for the attack.
@@
    AA-ADAdaptive Attack / Adaptive Defense: The exact setup as scenario 7, but the best-case for the defense is reported instead.
In each successive scenario, we evaluate using the most successful past policy, following standard practice in Markov Games: the worst-case opponent policy is fixed, and a best response to it is learned [32, 39]. Fixing other policies when computing a best response stabilizes learning in multi-agent environments, as it simplifies the problem to a single-agent setting – one that, as discussed in Section III-C, can be solved with standard RL.
Comparison to SotA. In Scenarios 0-8 we evaluate all possible combinations between (adaptive) attack and defenses. As a baseline to compare to, we additionally evaluate our approach to the state-of-the-art stateful defenses and adaptive attacks, that is Blacklight [19] and OARS [18] respectively. We implement both Blacklight and OARS in our interactive environments by using their publicly available code and parameters. As our environments do not return a rejection signal and to make a fair comparison, for OARS rejection coincides with a non-adversarial decision. We thus define 5 further scenarios:
@@
    VA-BDVanilla Attack / Blacklight Defense: Baseline performance of the attacks against Blacklight.
@@
    OA-BDOARS Attack / Blacklight Defense: OARS against Blacklight.
@@
    AA-BDAdaptive Attack / Blacklight Defense: Our adaptive attack against Blacklight.
@@
Our experiments were run on multiple machines, yet to give an idea for the time complexity of our defense, on an Intel i7-7700 CPU one forward pass in CIFAR – that is one response to one query – takes $8\pm 1.4$ ms for $\sim$700 MFLOPs.
@@ V-B Results
For consistency and comparability between evaluations, all results are from the same 100 test episodes. The gap value denotes the $L_{2}$ perturbation that initially separates the starting and the original samples, averaged over the 100 episodes. By testing the trained agents on budgets higher than 5K we discovered that the trend in reducing $L_{2}$ holds; to make the agent training tractable and the evaluation wider however, we limit the maximum query budget per adversarial episode to 5K.  Table II & Table III report the results for CIFAR10, while  Table IV reports MNIST. The closer examination of the empirical results help us extract and highlight several important insights, practical observations, and general implications for the broader AML field:
@@
  • The performance of both attacks deteriorates considerably against active defenses (VD), however the defenses reach their full potential only when also adaptive (AD).
  • Our adaptive defense (AD) outperforms both Blacklight (BD) and non-adaptive (VD) defenses, also when transferred (cf. Sc.12). Compared to Blacklight, it reduces ASR by $\sim$90(in HSJA) when trained against OARS, while it offers similar protection when transferred from another attack.
  • Overall, against the strongest attacks and in the worst case (Sc.7) for it, our defense contains the ASR in the range of $1-36\%$.
  • Our adaptive attack (AA) outperforms OARS (Sc.10 & 11) and vanilla attacks (VA) by a wide margin, without access to rejection sampling, and irrespective of the defense it faces; the one exception is our adaptive defense (AD), against which it has very limited success.
  • The advantage of adaptive attacks is more pronounced against active defenses, where they significantly outperform non-adaptive versions, cf. Sc.0→1, Sc.2→3.
@@
  • Evasive transformations interfere with the attack policy, as illustrated by the difference between BAGS and HSJA in Sc.5. For an attack to reach its full potential, these two should be adaptively controlled together.
  • In Sc.1 to 8, agents train against the best opponent policy as previously discovered, and ASR oscillates since following a fixed policy enables the learning of an optimal counter to it. Over successive adaptations, ASR eventually plateaus, indicating an equilibrium for each specific dataset and attack combination. This is illustrated in Figure 4.
  • The first time active defenses resisted adaptive attacks was in Sc.5 of CIFAR; we employed evasive transformations from then onward.
  • Different attack fundamentals respond differently to active defenses; the gradient estimation stage of HSJA has a disadvantage against similarity detection, while the jump and binary stages have an advantage.
  • For HSJA, engineering a state the adaptive defense could learn on, merely by leveraging our knowledge of the attack and its geometric functioning, proved impossible. What did prove effective however, was pure computation [^2] This is reminiscent of Sutton’s Bitter Lesson [51], the observation that progress in AI is often driven by gains in computation rather than problem-specific expert knowledge.: we used Contrastive Learning [52] to learn an embedding from raw queries, then used as state that transfers exceedingly well to other attacks like BAGS.
@@ VI Discussion
Limitations. To keep the amount of evaluations practical, we narrowed the scope to targeted attacks and to $L_{2}$ as the more suitable norm for visual similarity. Targeted attacks are strictly more difficult to perform than untargeted, while for binary classification targeted and untargeted coincide; our framework, however, can accommodate any adversarial goal or metric. Another simplifying assumption we make is that only one attack can take place at a time; however, the queuing technique we use for incoming queries is readily extensible to handle concurrent attacks. While we demonstrate that our active defense does transfer between attacks, another possibility to explore is training the defense on queries from different kinds of attacks. Finally, in our evaluation we focus on a wide range of adaptive and non-adaptive scenarios where agents learn and adapt interactively, thus limiting the number of datasets we experiment with; we back our empirical study however with an extensive theoretical analysis that supports the generality of our findings independent of context.
Future Work. The AMG framework we introduce is general by design and can accommodate the learning of optimal offensive and defensive policies in any domain of interest beyond image classification. A promising path for future research is the extension of our adaptive attacks and defenses to other domains and modalities, for instance malware, bot, and network intrusion detection. This is specifically because our approach circumvents the main obstacle of mapping gradient-based perturbations to feasible objects (eg. binaries) and instead can function directly in the problem space [37]. Another compelling and formidable challenge is automating the adaptive evaluations in AML, that is adapting beyond a specification by inventing tools to bypass defenses and thus imparting controllability to adversarial tasks. Finally, in our work we considered opponent agency as part of the environment; other domains, like malware detection, might benefit from explicit opponent modeling.